How would you solve this? Mary is presently 12 years older than her sister. Four years ago she was twice as old. How old are they now?
Answer by stanbon(75887)

How would you solve this? Mary is presently 12 years older than her sister. Four years ago she was twice as old. How old are they now?
--------
Equations:
m = s + 12
m-4 = 2(s-4)
-------
Substitute for "m" and solve for "s":
s+12-4 = 2s-8
s = 16 yrs (sister's age now)
--------------
Solve for "m"::
m = s + 12
m = 16+12 = 28 yrs (Mary's age now)
-----------
Cheers,
Stan H.
-----------